# Global Dynamics of McKean-Vlasov SDEs via Stochastic Order

### v1

@@ Abstract
This paper studies the rich dynamics of general one-dimensional McKean-Vlasov stochastic differential equations based on the preservation of stochastic order. The existing work is limited to the case of additive noises, while our framework allows multiplicative noises. When the equation has multiple invariant measures, we show basins of attraction of certain invariant measures contain unbounded open sets in the $2$-Wasserstein space, which is vacant in previous research even for additive noises. Also, our main results target on total orderedness and finiteness of invariant measures, global convergence to order interval enclosed by two order-related invariant measures, alternating arrangement of invariant measures in terms of stability (locally attracting) and instability (connecting orbits). Our theorems cover a wide range of classical granular media equations, such as symmetric, flat-bottom and asymmetric double-well confinement potentials with quadratic interaction, multi-well landscapes, and double-well landscapes with perturbations. Specific values for the parameter ranges, explicit descriptions of attracting sets and phase diagrams are provided.
@@
Keywords: McKean-Vlasov SDEs, order-preserving semigroups, invariant measures, connecting orbits, basins of attraction, stochastic order
@@ 1 Introduction
This paper focuses on one-dimensional McKean-Vlasov processes governed by the following stochastic differential equation (SDE),
@@
where $\{B_{t}\}_{t\in\mathbb{R}}$ is a standard two-sided Brownian motion on $\mathbb{R}$, $\mathcal{L}(X_{t})$ is the law of the random variable $X_{t}$, $V:\mathbb{R}\to\mathbb{R}$ is a potential function, $W:\mathbb{R}\to\mathbb{R}$ is a local interaction, and $\sigma:\mathbb{R}\to\mathbb{R}$ is a diffusion term. The category of McKean-Vlasov SDEs refers to a class of SDEs whose coefficients depend on laws of solutions. McKean [52] introduced it as a generalization of Itô’s diffusion processes by considering a nonlinear version of Fokker-Planck equations. For (1.1), the corresponding Fokker-Planck partial differential equation (PDE) is given by
@@
where $\rho_{t}$ is the probability density function of $X_{t}$. After that, the aspect of particle approximation property was examined for McKean-Vlasov SDEs. The equation (1.1) can be deduced as the mean-filed limit (or hydrodynamic limit) of the following interacting particle system,
@@
where $\{B_{t}^{i}\}_{i=1}^{\infty}$ is a sequence of independent Brownian motions. As the particle number $N\to\infty$, a single particle $X_{t}^{i,N}$ or the empirical measure $\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}}$ converges to the continuum process $X_{t}$ in some sense. This is called “propagation of chaos”, and we refer to [21, 22, 53, 61] on this topic. Due to the origin from infinite-particle limits, McKean-Vlasov SDEs are natural mathematical models for large-scale phenomena. See [9] for galactic dynamics, [10] for plasma physics, and [59] for molecular and fluid dynamics. [11, 13, 45] use this kind of SDEs to model collective behaviour of different species. Another application on mathematical biology is [26, 28, 35, 49] to neuroscience. For optimal controls and differential games of McKean-Vlasov type, as well as study on machine learning and neural networks, see [15, 24, 70].
The equation (1.1) is a model for the time evolution of granular materials (see Benedetto-Caglioti-Pulvirenti [8]), so it is also called a granular media equation. A typical setting for the local interaction is attractive quadratic, namely
$W(x)=\frac{\theta}{2}x^{2},\quad\theta>0.$
As for the potential function, it has more choices ranging from double-well to multi-well shapes. Depending on the noise intensity and the number of critical points of the potential, the number of invariant measures of the equation (1.1) may vary. We refer to Dawson [27], Tugaut [66] for phase transitions on the number of invariant measures in double-well landscapes, and Alecio [1] for multi-well cases. Roughly speaking, if $V^{\prime}$ has $(2n-1)$ simple zeros, the McKean-Vlasov SDE (1.1) admits a unique invariant measure when $\sigma>\overline{\sigma}_{c}$ above some upper-critical value, and it admits $(2n-1)$ invariant measures when $\sigma<\underline{\sigma}_{c}$ below a lower-critical value. For more related work on phase transitions, see also [16, 29, 30].
Besides phase transitions on the number of invariant measures, convergence to invariant measures has been an active research direction. The study for the dynamics of (1.1) can be divided into two different cases according to uniqueness and non-uniqueness of invariant measures. If there is a unique invariant measure for (1.1), since the interacting particle system (1.3) also has a unique invariant measure, uniform-in-time propagation of chaos can give the convergence result. Malrieu [51] goes through this way for uniformly convex interaction terms, and obtains exponential convergence rate via a log-Sobolev inequality. Cattiaux-Guillin-Malrieu [20] relaxes to the non-uniformly convex case, and instead uses the method of synchronous couplings. We refer to [32, 46] and literature therein for more on uniform propagation of chaos, and [38] for uniform log-Sobolev inequalities of mean-filed particle systems.
Investigation of the associated PDE (1.2) is an alternative path to prove convergence to the unique invariant measure for (1.1). From this perspective, [5, 6] shows convergence to the unique invariant measure with vanishing potential ($V=0$) by Bakry-Émery’s $\Gamma$-calculus, while Benedetto-Caglioti-Carrillo-Pulvirenti [7] utilizes free energy as a Lyapunov functional when $V$ is convex. Actually, when $\sigma$ is a constant, the PDE (1.2) is a Wasserstein gradient flow of the free energy (see Ambrosio-Gigli-Savare [3]), which is a sum of entropy, potential energy and interaction energy,
$\mathcal{F}(\rho):=\frac{\sigma^{2}}{2}\int_{\mathbb{R}}\rho(x)\log\rho(x)\mathop{}\!\mathrm{d}x+\int_{\mathbb{R}}V(x)\rho(x)\mathop{}\!\mathrm{d}x+\frac{1}{2}\int_{\mathbb{R}\times\mathbb{R}}W(x-y)\rho(x)\rho(y)\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}y.$
By the “Riemannian structure” on Wasserstein spaces, also known as Otto’s calculus (see Otto [58]), free energy attains its global minimum at the unique invariant measure. Then nonlinear log-Sobolev inequalities, that link free energy and its dissipation, guarantee the exponential convergence. Carrillo-McCann-Villani [18] achieves this by checking Bakry-Émery criterion, and Carrillo-McCann-Villani [19] develops an abstract theory of gradient flows in length spaces to handle degenerate convexity. For other relevant work, see [4, 31] and refenrence therein for Bakry-Émery analysis, [23, 50] for log-Sobolev inequalties, and [17, 64] for entropy dissipation.
In the case of non-unique invariant measures, the dynamics is not as intuitive as the case of unique invariant measures. When the equation (1.1) has three invariant measures with a double-well potential, one of the predictions is local convergence. Now, propagation of chaos is hard to resort to, as the particle system (1.3) still has a unique invariant measure. Nonetheless, the gradient flow (1.2) remains workable, where free energy has two local minima and one local maximum, attained at three invariant measures respectively. Assuming qualitative smallness of the noise intensity $\sigma$, Tugaut [67] uses WJ-inequality to show the local convergence. See Bolley-Gentil-Guillin [12] for the derivation of WJ-inequalities and consequent spectral gaps in Wasserstein distance. Meanwhile, Monmarché-Reygner [56] executes a local version of log-Sobolev argument to obtain the local convergence in a sharp range of $\sigma$ for double-well landscapes. We refer to Tamura [62, 63] for early works using PDE techniques, and Cormier [25] by Lions derivatives. As for other directions than local convergence for the double-well case, the convergence to one of invariant measures is carried out by Tugaut [65], and on the other hand, Liu-Qu-Yao-Zhi [48] indicates that the central invariant measure is unstable and has connecting orbits to the other two.
The present paper is a continuation of [48], which establishes a general framework to study unstable dynamics in multi-dimensional cooperative McKean-Vlasov SDEs. In this paper, we investigate the richer dynamical behaviour of the one-dimensional equation (1.1) with non-unique invariant measures, from the stochastic-order-preserving perspective. For general case, Theorem 2.1 and Corollary 2.2 show the set of invariant measures is totally ordered and finite, and its infimum and supremum enclose a $\mathcal{P}_{2}$-compact order interval $[\underline{\nu},\overline{\nu}]$, which attracts any bounded set in the $2$-Wasserstein space $\mathcal{P}_{2}(\mathbb{R})$. The infimum $\underline{\nu}$ attracts all probability measures smaller than $\underline{\nu}$, and the supremum $\overline{\nu}$ attracts those greater than $\overline{\nu}$. Given the zero-crossing number of the derivative of the potential function, Theorem 2.3 and Corollary 2.4 unveil an alternating arrangement of invariant measures in terms of stability and instability, with stable ones $\underline{\nu}$, $\overline{\nu}$ at the beginning and the end. A stable invariant measure attracts an open neighbourhood, and an unstable invariant measure means there exists a connecting orbit away from it. We can further prove the infimum $\underline{\nu}$ and the supremum $\overline{\nu}$ attract unbounded open neighbourhoods in $\mathcal{P}_{2}(\mathbb{R})$. Additionally, Theorem 2.5-2.9 provide some applications of the main results to classical examples, and their phase diagrams are presented in Figure 1, 2.
@@
Our approach is based on order-preservation of the semigroup generated by the equation (1.1) on the $2$-Wasserstein space $\mathcal{P}_{2}(\mathbb{R})$. To be precise, take an attractive quadratic interaction, and let $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ be the semigroup of (1.1). Consider the stochastic order “$\leq_{\mathrm{st}}$” defined by
$\mu\leq_{\mathrm{st}}\nu\text{ if and only if }\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\mu\leq\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\nu\ \ \text{for all bounded increasing functions $f\colon\mathbb{R}\to\mathbb{R}$}.$
Then $P_{t}^{*}$ is order-preserving on $\mathcal{P}_{2}(\mathbb{R})$ (see Theorem 4.4), that is
$\mu\leq_{\mathrm{st}}\nu\ \text{ implies }\ P_{t}^{*}\mu\leq_{\mathrm{st}}P_{t}^{*}\nu\ \text{ for $t\geq 0$}.$
Generally speaking, our method is parallel to the gradient flow approach. When the invariant measure is unique, to show global convergence, the gradient flow approach relies on the decreasing free energy along each single orbit, and a functional inequality involving free energy and its dissipation. By contrast, we develop the global attractor structures of order-preserving semigroups based on the characterization of order-boundedness and the coming down from infinity property. For the case of multiple invariant measures, the gradient flow approach needs a local functional inequality to show local convergence, while we establish the local attractor structures by tracing solutions backward along connecting orbits. To the best of our knowledge, the gradient flow approach concerning the basin of attraction of a single invariant measure is limited to the case of additive noises and local convergence, when invariant measures are non-unique. The main advantage of the order-preserving method is that we step further to give unbounded basins of attraction, even for multiplicative noises.
To illustrate that the basins of attraction of $\underline{\nu}$ (infimum invariant measure), $\overline{\nu}$ (supremum invariant measure) in Theorem 2.1 are unbounded, we exhibit inside a class of measures with explicit densities. Precisely, there are initial conditions arbitrarily far from $\underline{\nu}$ (resp. $\overline{\nu}$), for which the solutions converge to $\underline{\nu}$ (resp. $\overline{\nu}$). Consider a family of probability measures $\{\mu_{m}\}_{m\in\mathbb{R}}$ on $\mathbb{R}$, whose densities are given by
From a calculation of Fokker-Planck equations, the invariant measures of the equation (1.1) are those $\mu_{m}$ satisfying the following self-consistency equation,
$\int_{\mathbb{R}}(x-m)\mathop{}\!\mathrm{d}\mu_{m}(x)=0.$
Since $\mu_{m_{1}}\leq_{\mathrm{st}}\mu_{m_{2}}$ for $m_{1}\leq m_{2}$ (see Remark 4.3), it means $\underline{\nu}=\mu_{\underline{m}}$ and $\overline{\nu}=\mu_{\overline{m}}$ for some $\underline{m}\leq\overline{m}$. In this case, the solution starting from $\mu_{m}$ converges to $\underline{\nu}$ if $m\leq\underline{m}$, and converges to $\overline{\nu}$ if $m\geq\overline{m}$. Corollary 2.4 also gives a complete partition on limiting behaviour of solutions starting from the totally ordered arc $\{\mu_{m}\}_{m\in\mathbb{R}}$ in (1.4).
Our method can even show the basins of attraction contain unbounded open sets in $\mathcal{P}_{2}(\mathbb{R})$. By translating probability measures, we obtain order related pairs under the stochastic order. Given a local neighbourhood attracted by the infimum invariant measure, its basin of attraction actually contains all translated copies of the neighbourhood along the descending direction of the stochastic order. Correspondingly, the basin of attraction of the supremum invariant measure contains every ascending copy of its attracting neighbourhood. It is worth mentioning the stochastic order is inconsistent with the Wasserstein topology. In fact, any order interval has empty interior in $\mathcal{P}_{2}(\mathbb{R})$ (see Proposition A.1), and resembles a “thin slice” as shown in Figure 1, 2. So attracting open sets by the method of preserving such an unsatisfactory partial order relation is counter-intuitive.
As another necessary ingredient in our approach, we give the maximal number of invariant measures (see Proposition 5.3). Due to the analyticity and convolution form of the self-consistency function, we show the number of invariant measures of the equation (1.1) is finite and at most the zero-crossing number of the potential derivative $V^{\prime}$, regardless of the ranges of $\theta$ and $\sigma$. This helps to make our main theorems quantitative, and overcomes difficulties caused by qualitative results and missing parameter in the previous literature on phase transitions (see e.g., [1, 66]). One the other hand, this proposition allows $V^{\prime}$ vanishes on some interval, which is out of the settings in [67]. Also, it provides an insight that the essential factor in the maximal number of invariant measures is the zero-crossing number of $V^{\prime}$, instead of the number of simple zeros of $V^{\prime}$.
To demonstrate the versatility of our theory, we apply our main theorems to the following examples, including symmetric, flat-bottom, asymmetric double-well potentials, multi-well potentials, and double-well with perturbations (see Theorem 2.5-2.9). Take $W(x)=\frac{\theta}{2}x^{2}$ with $\theta>0$, and take the potential derivatives respectively as
@@
[…]
In all these equations, all the dynamical results from Theorem 2.1, Corollary 2.2, Theorem 2.3 and Corollary 2.4 hold. The specific values for the ranges of $\theta$, $\sigma$, and explicit descriptions of attracting sets are given. For the phase diagrams, see Figure 1, 2.
This paper is organized as follows. In Section 2, we state our main results. Section 3 is about properties of the stochastic order, including an equivalent characterization of order boundedness in $\mathcal{P}_{2}(\mathbb{R})$ and a generalization of the lattice property. Section 4 is devoted to the proof of Theorem 2.1 and Corollary 2.2. We obtain the structure of global attractors of order-preserving semigroups, the comparison theorem for extrinsic-law-dependent SDEs, and the coming down from infinity property of McKean-Vlasov SDEs. Section 5 serves the proof of Theorem 2.3 and Corollary 2.4. We deduce the information on the number of invariant measures, the existence of positively invariant neighbourhoods of invariant measures, and the structure of local attractors of order-preserving semigroups. Also, the order bounds of limit sets are found by tracing solutions backward along connecting orbits. In Section 6, we apply the main theorems to examples stated in Theorem 2.5-2.9. In Appendix A, we show that, in $\mathcal{P}_{2}(\mathbb{R})$, any probability measure has no order relation with any open set, which denies the existence of strong ordered pairs for the stochastic order.
@@
[…]
[type=symbols,style=mylong]
2 Assumptions and Main Results

### v2

# Global Dynamics of Granular Media Equations via Stochastic Order

@@ Abstract
This paper studies the rich dynamics of one-dimensional granular media equations with attractive quadratic interactions. Building on the monotone dynamical systems framework developed in an earlier work, we allow for multiplicative noise, in contrast to most existing results restricted to additive noise. Within this framework, we show that, in the one-dimensional setting, invariant measures are totally ordered with respect to the stochastic order. The basins of attraction of the minimal and maximal invariant measures contain unbounded open sets in the $2$-Wasserstein space, which is vacant in previous research even for additive noises. Also, our main results address the global convergence to the order interval enclosed by the minimal and maximal invariant measures, and an alternating arrangement of invariant measures in terms of stability (locally attracting) and instability (as the backward limit of a connecting orbit). Our theorems cover a wide range of classical granular media equations, such as double-well and multi-well landscapes. Specific values for the parameter ranges, explicit descriptions of attracting sets and phase diagrams are provided.
@@
Keywords: Granular media equations, long-time behavior, invariant measures, basins of attraction, order-preserving semiflow, stochastic order
@@ 1 Introduction
The category of McKean-Vlasov SDEs refers to a class of SDEs whose coefficients depend on laws of solutions. McKean [38] introduced it as a generalization of Itô’s diffusion processes by considering a nonlinear version of Fokker-Planck equations. Among McKean–Vlasov SDEs, one of the most important and widely studied classes with concrete applications is given by granular media models:
@@
where $\{B_{t}\}_{t\in\mathbb{R}}$ is a standard two-sided Brownian motion on $\mathbb{R}$, $\mathcal{L}(X_{t})$ is the law of the random variable $X_{t}$, $V:\mathbb{R}\to\mathbb{R}$ is a confining potential function, $W:\mathbb{R}\to\mathbb{R}$ is an interaction potential, $\sigma:\mathbb{R}\to\mathbb{R}$ is a diffusion term and
$(W^{\prime}*\mu)(x):=\int_{\mathbb{R}}W^{\prime}(x-y)\mathop{}\!\mathrm{d}\mu(y),\ \text{ for all probability measure $\mu$ on $\mathbb{R}$.}$
The law $\rho_{t}=\mathcal{L}(X_{t})$ of the solution to (1.1) formally solves the associated nonlinear Fokker-Planck equation
@@
which describes the evolution of the particle density. From another point of view, The equation (1.1) can be deduced as the mean-field limit (or hydrodynamic limit) of the following interacting particle system,
@@
where $\{B_{t}^{i}\}_{i=1}^{\infty}$ is a sequence of independent Brownian motions. As the particle number $N\to\infty$, a single particle $X_{t}^{i,N}$ or the empirical measure $\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}}$ converges to the continuum process $X_{t}$ in some sense. This is called “propagation of chaos”, and we refer to [10, 11, 39, 48] on this topic. Due to their origin as infinite-particle limits, equations (1.1)-(1.2) provide natural mathematical models for large-scale phenomena. They are widely applied in various fields, such as battery models [21], optimization [6] and machine learning [40].
In this paper, we focus on the typical attractive quadratic interaction potential, namely
$W(x)=\frac{\theta}{2}x^{2},\quad\theta>0,$
which arises naturally in a variety of application contexts, including muscle contraction [33], chemical kinetics [29], statistical physics [23], and large economic systems [3]. A salient feature of this class of equations is the occurrence of multiple invariant measures if the noise is not too strong. For the double-well landscapes with additive noise, Dawson [13] and Tugaut [52] obtain phase transitions in the number of invariant measures. For the multi-well case, we refer to Alecio [1] and the references therein. Roughly speaking, this phenomenon can be understood through the geometry of the confining potential. It is common to consider well-shaped confining potentials $V$ in granular media equations, which means $V(x)\to\infty$ as $x\to\pm\infty$. If its derivative $V^{\prime}$ only has simple zeros, then basically $V^{\prime}$ has odd (2n1)(2n-1) zeros. In such a setting, the equation has exactly (2n1)(2n-1) invariant measures when noise is small enough, and has a unique invariant measure when noise is large.
Consequently, the study of the dynamics of (1.1) is often divided into two regimes: the case of a unique invariant measure and the case of multiple invariant measures. In the former, convergence to the unique invariant measure has been extensively studied. For instance, Carrillo-McCann-Villani [7] employed generalized logarithmic Sobolev inequalities and mass transportation inequalities, while Cattiaux-Guillin-Malrieu [9] used a uniform-in-time propagation of chaos method to demonstrate convergence to the unique invariant measure. For more on this topic, see [8, 16, 22, 56] and the references therein.
In the presence of multiple invariant measures, a complete characterization of the basins of attraction is in general a highly challenging problem. Current results in this direction are mainly restricted to establishing the local convergence around certain stable invariant measures, showing that solutions starting from initial distributions sufficiently close to them converge to the corresponding invariant measures. In Tugaut [53], local convergence is established under the assumptions that the initial condition admits a finite higher-order moment, a $C^{\infty}$-smooth density, and finite entropy. Monmarché-Reygner [43] applied a local version of the log-Sobolev inequality to obtain local convergence within a sharp range of the noise intensity $\sigma$ for double-well landscapes. Zhang [58] achieved local convergence by linearizing the nonlinear Markov semigroup, applicable to double-well landscapes. For early works using the free energy function, see Tamura [49, 50], while approaches relying on Lions derivatives can be found in Cormier [12]. From a global perspective, Tugaut [51] showed that, for double-well granular media equations with qualitatively sufficiently small noise, solutions starting from initial conditions with finite higher-order moment, a $C^{\infty}$-smooth density, and finite entropy will converge to one of the three invariant measures.
In [36], a monotone dynamical systems framework for cooperative McKean–Vlasov SDEs was developed. More precisely, under a local dissipativity assumption, the existence of multiple order related invariant measures in the Wasserstein space was established, together with monotone connecting (heteroclinic) orbits between them with respect to the stochastic order.
Building on the main result of [36], the present paper goes further in the one-dimensional setting. By analyzing the local dissipativity condition and the number of invariant measures, we identify connecting orbits for a broader class of one-dimensional granular media equations with multiplicative noise, including, in particular, polynomial-type confining potentials of even degree. Relying on the existence of such connecting orbits together with a structural theorem for local attractors of order-preserving semigroups, we prove local convergence in the $2$-Wasserstein space. Furthermore, by combining a structural result on global attractors with an analysis of the stochastic order, we show that the basins of attraction of the minimal and maximal invariant measures contain unbounded open sets (see our main results Theorem 1.2-Corollary 1.5). To illustrate the applicability of our abstract results, we present below a theorem for the case where $V^{\prime}$ has finitely many simple zeros.
Theorem 1.1.
Suppose Assumption 1 holds, and $V^{\prime}$ has finite zeros and they are all simple. Then there exist $\theta_{0}>0,\ \sigma_{0}>0$ such that for all $\theta\geq\theta_{0}$ and $\overline{\sigma}\leq\sigma_{0}$,
  1. (i)
    (Total Orderedness) there are exactly (2n1)(2n-1) invariant measures $\{\nu_{k}\}_{k=1}^{2n-1}$ with $\nu_{1}<_{\mathrm{st}}\nu_{2}<_{\mathrm{st}}\cdots<_{\mathrm{st}}\nu_{2n-1}$ for some integer $n\geq 1$;
  2. (ii)
  3. (iii)
    (Locally Attracting) there exist $r>0$ and $a_{1}<\cdots<a_{k}$ such that for all $1\leq k\leq n$, $\nu_{2k-1}\in B(\delta_{a_{k}},r)$ and
    $\mathcal{W}_{2}(P_{t}^{*}B(\delta_{a_{k}},r),\nu_{2k-1})=\sup_{\mu\in B(\delta_{a_{k}},r)}\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2k-1})\to 0\text{ as }t\to\infty;$
  4. (iv)
    (Attracting Unbounded Open Neighbourhoods) if $\mu\leq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{1}},r)$, then
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0\text{ as }t\to\infty.$
    If $\mu\geq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{n}},r)$, then
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2n-1})\to 0\text{ as }t\to\infty.$
    In particular,
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0\text{ as }t\to\infty\ \text{ for all $\mu\in\bigcup_{a\leq 0}B(\delta_{a_{1}+a},r)$},$
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2n-1})\to 0\text{ as }t\to\infty\ \text{ for all $\mu\in\bigcup_{a\geq 0}B(\delta_{a_{n}+a},r)$}.$
Our theorem clearly includes confining potentials of polynomial type, in particular those with
$V^{\prime}(x)=a\prod_{k=1}^{2n-1}(x-\alpha_{k}),\qquad a>0,\ \alpha_{k}\in\mathbb{R}\ \text{and}\ \alpha_{i}\neq\alpha_{j}\ (i\neq j).$
We stress that our approach is expected to extend to more general interaction potentials, for instance those satisfying $W^{\prime\prime}>0$, as well as to higher-dimensional systems, including a multi-species population models; see [14, 15].
@@
Our approach begins with the order-preserving property of the law-evolution semigroup generated by equation (1.1) on the $2$-Wasserstein space $\mathcal{P}_{2}(\mathbb{R})$ (see e.g., [30, 36, 45]). More precisely, we consider the stochastic order $\leq_{\mathrm{st}}$ defined by
@@
Under a hyper-dissipativity assumption, we develop a dynamical result describing the global attractor structure of order-preserving semigroups (Proposition 3.2), based on a characterization of order-boundedness (Proposition 2.2) and the coming down from infinity property (Proposition 3.6). This yields a partial description of basins of attraction in the sense of the stochastic order (Theorem 1.2, Corollary 1.3). Building upon this global attractor analysis, and under the additional assumptions on local dissipativity and zero-crossing number, we further derive local attractor structures by tracing solutions backward along connecting orbits (Section 4.2). As a consequence, we obtain local convergence in the topological sense, as well as the existence of unbounded open subsets contained in the corresponding basins of attraction (Theorem 1.4, Corollary 1.5). It is worth noting that this phenomenon is far from obvious. Indeed, any order interval has empty interior in $\mathcal{P}_{2}(\mathbb{R})$ (see Proposition A.1), and resembles a ‘thin slice” as shown in Figure 1, 2. So attracting open sets by the method of preserving such an unsatisfactory partial order relation is counter-intuitive.
Finally, our results suggest the presence of a saddle-point structure in the double-well granular media equation. This leads us to formulate, at the end of the paper, a conjecture on the existence of an unordered invariant separatrix containing the unstable invariant measure, which separates the basins of attraction of another two stable invariant measures.
This paper is organized as follows. In Section 1.1, we state the main results beyond Theorem 1.1. Section 2 is about properties of the stochastic order, including an equivalent characterization of order boundedness in $\mathcal{P}_{2}(\mathbb{R})$ and a generalization of the lattice property. Section 3 is devoted to the proof of Theorem 1.2 and Corollary 1.3. Section 4 serves the proof of the remaining theorems and the formulation of a saddle-point conjecture in the double-well granular media equation. In Appendix A, we show that, in $\mathcal{P}_{2}(\mathbb{R})$, any probability measure has no order relation with any open set, which denies the existence of strong ordered pairs for the stochastic order.
@@
1.1 Assumptions and Main Results